### v2

@@ 3 Discussions
Our M3FM model is the first-of-its-kind medical multimodal multitask foundation model that has been successfully aligned with the clinical practice with an initial emphasis on chest CT especially lung cancer screening. As the currently most promising AI technology, multimodal multitask foundation models are highly desired to integrate clinical workflows and improve their performance and efficiency. The progress along this direction has not been as fast as that with large language models or simple language-vision models because of the major data and model challenges in the medical domain. To overcome the data challenge, we follow the clinical practice and develop a complete workflow to curate medical multimodal multitask datasets, which include medial tasks of interest definition, task-specific multimodal data collection, multimodal data preprocessing and alignment, and multimodal question-answering dataset construction. Consequently, our constructed dataset is the first and largest medical multimodal multitask dataset and systematically aligned across 3D tomographic medical images, diverse clinical data, and LCS-related clinical tasks. Moreover, the presented data integration workflow can be easily applied to enlarge the existing datasets as it does not require extra labeling efforts. Modeling high-dimensional medical multimodal multitask datasets presents several unique challenges. These include encoding 3D tomographic images that feature abnormalities of varying sizes, effectively representing diverse clinical data types, seamlessly integrating multiple tasks, and efficiently handling computational costs. By innovatively addressing these specific challenges, we have prototyped a scalable and adaptable foundational model in a unified framework with a user-friendly multimodal question-answering interface. Our M3FM models were trained and evaluated on large and diverse datasets collected from multiple medical institutes. M3FM models significantly outperform the previous best task-specific models across multiple medical tasks and exhibit scalability to larger datasets and more medical tasks.
The M3FM developed and tested in our study has limitations and clinical implications. Our testing was retrospective and offline rather than in a prospective, real-world reporting and patient management environment. A clinical environment could pose logistical challenges in real-time data curation and display with information streams across multiple distinctive and yet inter-connected interfaces including electronic medical records, radiology information systems (dictation or reporting interface) and PACS. We did not test the clinical impact of M3FM either in radiology or post-radiology care scenarios. Likewise, we did not evaluate the best and targeted method of information display to improve decision-making with multimodal information without inundating physicians and compromising their workflow efficiency.
Within the scope of these limitations, there are opportunities for clinical impacts of M3FM. With regulatory clearance and integration with real-time clinical workflows, M3FM can provide a real-time, dynamic, and customizable dashboard for information summary and decision-making. For example, during reporting of lung findings, it would be helpful to have a display of M3FM-derived results, previously reported lung findings and lung-specific clinical findings (such as COPD, prior lung nodules, and LungRADS categories), and during cardiovascular “field” reporting, to have a display of M3FM-based estimates, cardiac risks factors from past medical history and laboratory/ECG/echocardiography reports. Beyond LCS, such a dashboard can help in various other ways; for example, in cancer staging, we can access not only M3FM outputs but also medical, radiation, and surgical treatments at the time of determining the best course of management.
M3FM models set the new performance records on all 17 tasks of interest on the OpenM3Chest datasets. Compared to existing models that focused on single-modality, task-specific applications, M3FM achieved a substantial improvement in accuracy. These results on clinical datasets highlight the potential of medical foundation models for precise medical AI. Specifically, our experimental results affirm that larger-scale models yield better outcomes in medical multimodal multitask performance. Nevertheless, the performance gain when upgrading from M3FM-Large to M3FM-Huge is less pronounced than that when upgrading from M3FM-Base to M3FM-Large. We believe that the limited improvement with the huge model could be substantially attributed to the size and quality of the current datasets, and with even larger data size and finer data quality, we expect to have even better performance, given the complexity of the underlying clinical tasks and by the scaling laws demonstrated in the NLP domain [38]. Hence, our findings highlight the necessity to gather, curate and integrate extensive, high-quality, multimodal, multitask datasets.
@@
M3FM models effectively encode multimodal data and flexibly synergize different medical tasks. Importantly, different tasks may have different combinations of multimodal inputs and varying output formats. Generally, our findings indicate that multimodal modeling can enhance performance on specific tasks, while multitask learning tends to improve outcomes across a broader range of tasks. Still, some tasks may only depend on single modality data; e.g., lung nodule detection and characterization do not benefit from clinical datasets more than LDCT scans in our experiments. The overall results from multimodal and multitask modeling efforts are encouraging, indicating that M3FM possesses the potential to encompass a wider array of medical tasks and datasets for unified and improved medical AI applications.
@@
M3FM models show better generalizability to independent datasets than task-specific models. Our experimental results show that the multimodal multitask M3FM models achieved consistently and significantly better results by up to 20+than the previous models trained in the same settings. However, for some tasks, the multi-modality modeling may decrease the generalizability relative to the single-modality modeling. This could be due to the variability in data collection procedures and standards. Thus, it is important to design a standardized and robust data resourcing and collecting pipeline. In all the experiments, multitask learning can consistently improve generalizability.
@@
M3FM models are capable of identifying informative clinical elements both quantitatively and qualitatively. Specifically, our M3FM models have uncovered a strong correlation between CVD diagnosis, CVD mortality prediction, and the historical presence of heart disease/attacks, hypertension, stroke, and diabetes through ablation inference. This suggests that altering the combination of clinical data inputs fed into the trained model affects the quantitative outcomes on evaluation datasets, thereby pinpointing the informative clinical elements. Quantitatively, attention maps can be visualized for both image and text inputs, illuminating the elements that correlate with predictions. This not only showcases the interpretability of M3FM models but also aids in understanding the relationships between clinical data and diseases, facilitating optimization of multimodal multitask performance.
@@
M3FM models significantly improve multimodal modeling for out-of-distribution tasks. A key feature of foundation models is their ability to aid tasks beyond those defined by the training datasets. In this study, we finetuned our M3FM models for immunotherapy prognosis prediction, an out-of-distribution task characterized by entirely different clinical inputs. Our experiments demonstrated that the pre-trained M3FM model substantially enhances the new task performance on a relatively small dataset. This capability is particularly valuable when expanding some clinical datasets is challenging due to data rarity and associated cost.
@@
The exceptional performance of M3FM models heralds a promising avenue for leveraging multimodal data and multiple tasks in developing AI-empowered, cost-effective, and superior healthcare solutions. Within the scope of lung cancer screening (LCS) in particular, we plan to create a more robust M3FM model that can broaden and refine LCS implementation, ultimately reducing lung cancer mortality. Beyond LCS, we are actively work for M3FMs to be translated on a comprehensive medical multimodal multitask platform. This M3FM system would rapidly evolve and accommodate increasingly more medical tasks with diverse multimodal data combinations in the form of specialized and generalized medical AI models.
Figure 9: General Workflow for Medical Multimodal Multitask Dataset Construction.
@@ 4 Methods
4.1 Dataset Construction Procedure
Figure 9 presents our general workflow for constructing the multimodal multitask medical datasets. First, we define a set of medical tasks for specific clinical scenarios and collect the relevant multimodal datasets for these tasks from various medical sources. Then, we preprocess the multimodal datasets to accumulate information such as the organ of interest localization in 3D CT images, key phrases in radiology reports, and ground-truth labels from different resources, and align all multimodal data for each patient. Subsequently, we create questions and answers for each specific task according to the clinical practice. Note that there is no need for extra labeling efforts by radiologists across the whole workflow so that large-scale medical multimodal multitask datasets can be systematically and cost-effectively constructed.
Figure 10: Technical details of the M3FM model in Figure 1: (a) the overall M3FM architecture with CTViT, text transformers, associated encoders, and predictors; (b) image encoder; (c) text encoder; and (d) task encoder.
@@ 4.2.1 Overall architecture
Our medical multimodal multitask foundation model can effectively encode multimodal data and flexibly perform multitasks via text prompting in a unified and scalable fashion. As shown in Figure 1, M3FM consists of the four main components: CTViT, text Transformer, task encoder, and predictors. The key details of each component are given in Figure 10. CTViT takes volumetric CT images of varying sizes as inputs, extracts multi-scale image patches from them, and computes discriminative features of these patches. The text Transformer produces the embeddings of clinical text and the embeddings of textual questions respectively. Given any combination of image, text, and task token embeddings, the task encoder extracts the task-specific features corresponding to the special $<TASK>$ token. Finally, the task-specific predictor outputs the final answer from the task-specific features of the integrated multimodal data. In the following, we will describe each of these components in detail.
@@ 4.2.2 CTViT
CTViT extracts embedding features of multi-scale 3D CT volumes with physical size awareness. CTViT has two parts: a multi-scale CT tokenizer and an image encoder. To process a 3D CT scan, we divide each image volume into non-overlapped 3D patches as in [56]. Each 3D patch is referred to as an image token. Since various diseases are at different scales in the CT images, we design a multi-scale CT tokenizer, which consists of multiple linear embedding layers corresponding to different sizes of image patches, as shown in Figure 10(a). Each embedding layer has a linear transformation and a set of learnable positional embeddings. Each image token embedding is the sum of its linear transformation and the positional embedding. All sizes of image tokens are mapped to the same image embedding space. Inspired by [54], we decompose the 3D position embedding into two parts indexing in-plane and through-plane positions respectively. In other words, we have two positional embeddings: one for the 2D space within each slice and the other for the 1D range of slice position. The 3D positional embedding is the sum of them. By doing so, the number of learned parameters can be reduced. Figure 10(b) shows the image encoder in detail. Different CT scans may have different physical sizes specific to the individual patient size. The physical size is an important factor in some clinical tasks. Thus, we encode the physical size for image tokens with sine-cosine functions of different frequencies and add it to the image token embedding. Then, the image encoder was implemented as the plain ViT [57] that consists of multiple self-attention Transformer layers and a subsequent linear transformation layer that maps the image embedding space to the task embedding space. By disentangling physical size from image contents, we can flexibly perceive any size of CT volumes with physical size awareness without resampling CT volumes to have a consistent image tensor across different inputs.
@@ 4.2.3 Text Transformer
Any decent language model can be utilized as the text Transformer in M3FM. Here the text Transformer was implemented as a Byte-level Byte-Pair-Encoding (BBPE) tokenizer [52], a text encoder consisting of the original Transformer layers [16], and a linear transformation layer, as shown in Figure 10(c). On one hand, the text encoder encodes patient-specific clinical information, such as demographics, smoking history, disease history, cancer history, and other clinical data, which are free text; e.g., “The patient is 56.0 years old. Gender is Female. Ethnicity is neither Hispanic nor Latino. Height is 60.0 inches. Weight is 105.0 pounds. Education is associate degree/ some college. Former smoker. Smoking duration is 38.0 pack years. Smoking intensity is 20.0 cigarettes per day. 2.0 years since quit smoking. The patient had asthma (childhood) diagnosed at 7.0 years old. The patient had hypertension diagnosed at 53.0 years old. The patient had pneumonia diagnosed at 50.0 years old. Patient’s brother(s) (including half-brothers) have lung cancer.” On the other hand, the text encoder encodes free-text task instructions/questions, which are used as the input of the task encoder to extract task-specific embedding features from the multimodal data; e.g., “Is there any significant cardiovascular abnormality?” and “Predict the lung cancer risk over 6 years.” This approach allows for embedding any combination of clinical information through free-text prompting, regardless of order. The control signals for specific tasks are then extracted from the text prompts by the same text encoder. Again, the linear transform maps the text embedding space to the unified task embedding space. For all our M3FM models, the number of Transformer layers is 12, and the size of text token embeddings is 768.

### v3

@@ 3 Discussions
The contributions of the proposed M3FM can be summarized in two main aspects. First, as the first-of-its-kind SOMAI model for LCS, M3FM effectively encodes multimodal medical data including arbitrary combinations of multi-scale 3D tomographic images and various clinical data, and flexibly performs multiple tasks via free-text prompting. In particular, our CT Vision Transformer (CTViT) is a unique component designed to perceive 3D CT images. CTViT can flexibly process multiple image sizes through our multi-scale linear tokenizer and disentangled physical size embedding designs. We developed the corresponding self-supervised learning training algorithm that facilitates the pre-training of the multi-scale CTViT on large 3D CT datasets. To make M3FM scalable across multiple tasks, we designed a distributed task-parallel training strategy, which assigns a single task to each device while allowing different devices to process different inputs/outputs for multitask parallel optimization. Second, we presented the whole workflow for the SOGMAI model development, from clinical multitask definition to multimodal data curation, from radiologists’ reading procedure to the unified MQA framework, and from self-supervised pre-training to synergistic multitasking with high-dimensional multimodal data. In particular, our MQA framework is akin to how medical professionals perform multiple tasks while considering multimodal data, naturally allowing unified training and interactive inference. Importantly, the whole pipeline is designed with scalability, allowing M3FM to be readily scaled up by integrating more training datasets and undertaking a broader range of clinical tasks.
The M3FMs set the new performance records on all tasks of interest on our curated large-scale OpenM3Chest datasets collected from multiple medical institutes. The M3FMs significantly outperformed the previous models, which mainly focus on task-specific and single modality. Promisingly, our experimental results affirm that the larger M3FM produces better outcomes in medical multimodal multitask settings. These positive outcomes underscore the importance of systematically collecting and curating large-scale, multimodal, multitask datasets to enhance the precision of medical AI applications. All these findings on clinical datasets highlight the potential of SOGMAI.
@@
The M3FMs effectively encode multimodal data and flexibly synergize different medical tasks. Importantly, different tasks may have different combinations of multimodal inputs and varying output formats. Generally, our findings indicate that multimodal modeling can enhance performance on specific tasks, while multitask learning tends to improve outcomes across a broader range of tasks. Still, some tasks may only depend on single modality data; e.g., lung nodule detection and characterization do not benefit from clinical datasets more than LDCT scans in our experiments. The overall results from multimodal and multitask modeling efforts are encouraging, indicating that M3FM possesses the potential to encompass a wider array of medical tasks and datasets for unified and improved medical AI applications.
Nevertheless, there are certain limitations to current M3FM results. The above evaluation was retrospective and offline rather than in a prospective, real-world reporting and patient management environment. In current clinical systems, real-time multimodal data integration could pose logistical challenges, particularly when managing information streams across multiple, distinct yet interconnected interfaces such as electronic medical records, radiology information systems, and picture archiving and communication systems. We did not test the clinical impact of M3FM either in radiology or post-radiology care scenarios. Likewise, we did not evaluate the most effective method of information display to improve decision-making with multimodal information without inundating physicians and compromising their workflow efficiency. While our experimental results affirm that larger-scale models yield better outcomes in medical multimodal multitask performance, the performance gain when upgrading from M3FM-Large to M3FM-Huge is less impressive than when upgrading from M3FM-Base to M3FM-Large. We believe that this limited improvement could be substantially attributed to the size and quality of the current datasets, and with even larger and better datasets we expect to have superior performance, by the scaling laws. Within the scope of these limitations, there are opportunities for clinical impacts of M3FM. With regulatory clearance and integration with real-time clinical workflows, M3FM can provide a dynamic, and customizable dashboard for information summary and decision-making. For example, during the reporting of lung findings, it would be helpful to have a display of M3FM-derived results, previously reported lung findings, and lung-specific clinical findings (such as Chronic Obstructive Pulmonary Disease (COPD), prior lung nodules, and LungRADS categories), and during cardiovascular “field” reporting, to have a display of M3FM-based estimates, cardiac risks factors from past medical history and laboratory/ECG/echocardiography reports. Beyond LCS, such a dashboard can help in various other ways; for example, in cancer staging, we can access not only M3FM outputs but also medical, radiation, and surgical treatments when determining the best management.
As discussed in [43], there are specific concerns for AI in medicine, such as generalizability, explainability, adaptability, etc. This study has demonstrated initial efforts in addressing AI-specific concerns.
@@
The M3FMs show better generalizability to independent WFUSM and MGH datasets than task-specific models. It is worth mentioning that the independent evaluation is prospective in terms of the date of data collection. Our experimental results show that the multimodal multitask M3FM models achieved consistently and significantly better results by up to 20+than the previous models trained in the same settings. However, for some tasks, the multi-modality modeling may decrease the generalizability relative to the single-modality modeling. This could be due to the variability in data collection procedures and standards. Thus, it is important to design a standardized and robust data resourcing and collecting pipeline. In all the experiments, multitask learning can consistently improve generalizability.
@@
The M3FMs are capable of identifying informative clinical elements both quantitatively and qualitatively, which offers certain explainability. It is achieved with the MQA framework and the attention mechanism. Specifically, the MQA framework naturally allows users to examine the response changes to different combinations of imaging and clinical data, and thus, the informative clinical elements can be identified as those contributing to statistically high prediction accuracy. Our M3FMs have uncovered a strong positive correlation between CVD diagnosis, CVD mortality prediction, and the historical presence of heart disease/attacks, hypertension, stroke, and diabetes through ablation inference. Assuming this discovery is not a piece of common knowledge, M3FM would lead to new biomedical findings. Quantitatively, attention maps can be visualized for both image and text inputs through the attention mechanism, illuminating the elements that correlate with predictions. This visualization offers certain the interpretability of M3FM models in understanding the relationships between clinical data and diseases.
@@
The M3FM has the adaptability to significantly improve multimodal modeling for out-of-distribution tasks through transfer learning. A key feature of foundation models is their ability to aid tasks beyond those defined by the training datasets. In this study, we finetuned our M3FMs for immunotherapy prognosis prediction, an out-of-distribution task characterized by entirely different clinical inputs. Our experiments demonstrated that the pre-trained M3FM model substantially enhances the new task performance on a relatively small dataset. This capability is particularly valuable when expanding some clinical datasets is challenging due to data rarity and associated costs.
@@
In conclusion, the unified architecture and exceptional performance of the M3FMs herald a promising avenue for leveraging multimodal data and clinical multitasks in developing AI-empowered, specialty-oriented superior healthcare solutions. Within the scope of LCS in particular, we see the potential to translate the M3FM model on our collaborative clinical sites, broaden and refine LCS implementation, and ultimately reduce lung cancer mortality. Hopefully, our M3FM system would become an effective platform to accommodate more medical tasks with diverse multimodal data combinations, from specialized to increasingly more generalized medical AI models.
@@ 4 Methods
4.1 Medical Multimodal-Multitask Dataset Construction
Figure 2-(a) presents our general workflow for constructing the multimodal multitask medical datasets, which consists of four main steps: (1) medical tasks of interest definition; (2) task-specific multimodal data collection; (3) multimodal data processing and alignment; and (4) MQA construction. The details for the first two steps are described in Section 2.1. Here we introduce the third and fourth steps.
The multimodal data processing to select qualified multimodal data and prepare them for the alignment, including the CT data processing, clinical data processing, and ground-truth label calculating. In CT data processing, we localize the sub-volumes that mainly contain the task-relevant regions in the 3D CT using a segmentation model [44]. Specifically, we segment three parts, i.e., left lung, right lung, and heart regions consisting of the myocardium, left/right atrium, left/right ventricle, and pulmonary artery. It is worth underlining that the precision of the segmentation results does not need to be extremely high. Our primary objective is to utilize rectangular boxes to wrap the segmented areas, ensuring that task-relevant sub-volumes are included and extraneous regions are disregarded. We excluded the CT series having less than 64 axial slices in all collected datasets. For each CT series, the reconstruction voxel sizes are prepared in the axial, coronal, and sagittal dimensions, which will be used as input to the CTViT. The clinical data processing is to represent various combinations of clinical data within free text. We have established a specific sentence format for each clinical element, as detailed in Table S1 of the supplementary. The final free-text clinical data for each examination is constructed by aggregating the sentences corresponding to all available and positive clinical data, as illustrated by the text inputs in Figures 1 and 5. In ground-truth label calculating, we first extract task-specific label information from different sources (see Table 1 for specific label sources of each task) and then combine all information to calculate the label. The details for each task ground-truth label calculation are described in Table S2 of the supplementary. Next, we align the clinical data in free text, the CT data with segmented parts and physical size, and the calculated labels for all exams in each task. In particular, each task anatomically corresponds to the segmented CT sub-volume. The key principle here is to remove the irrelevant image regions to reduce the computation cost while keeping the original resolution without losing the information. The task-specific CT sub-volumes are illustrated in the image input column of Table 1. Specifically, 2.5D left or right lung sub-volume is used as the image input for lung nodule detection and characterization. We fixed the number of slices to 16 for each input considering that a lung nodule is usually tiny relative to the whole lungs. For the nodule-presented case, the location labels of the left/right lung, slice number, and the bounding box coordinates are used to crop the target sub-volume. For the non-presented nodule case, the input sub-volume is randomly cropped within the segmented lung regions. The CVD tasks use a 3D rectangle box wrapping all heart regions as the image input. For lung cancer risk prediction, we separately input the left or right lung in a 3D rectangle box. In the training datasets, the location of the left and/or right lung where lung cancer is presented is required to align each lung with the risk labels. In the validation and test datasets, the location of lung cancer is not required as the ground-truth labels are patient-level and the predicted lung cancer risks are the maximum scores between two lungs for each patient. For all other chest disease diagnosis tasks, a 3D rectangle box wrapping both lungs is used as the image input. Subsequently, the MQA construction is to create questions and answers for each specific task with the aligned multimodal data, and the resulting MQA datasets define the model’s input and output formats. In Table 1, one example question and the corresponding answer candidates are presented for each task. In the training stage, ten different questions for each task were used as shown in Table S3 of the supplementary. Note that there is no need for extra labeling efforts by radiologists across the whole workflow so that large-scale medical multimodal multitask datasets can be systematically and cost-effectively constructed.
Figure 6: Technical details of the M3FM in Figure 1: (a) the overall M3FM architecture with CTViT, text transformers, associated encoders, and predictors; (b) CTViT pretraining. (c) image encoder; (d) text encoder; and (e) task encoder.
@@ 4.2.1 Overall architecture
Our medical multimodal multitask foundation model is designed to effectively encode multimodal data and flexibly perform multitasks via text prompting in a unified and scalable fashion. As shown in Figure 1, M3FM consists of the four main components: CTViT, text Transformer, task encoder, and predictors. The key details of each component are given in Figure 6. CTViT takes volumetric CT images of varying sizes as inputs, extracts multi-scale image patches from them, and computes discriminative features of these patches. The text Transformer produces the embeddings of clinical text and the embeddings of textual questions respectively. Given any combination of image, text, and task token embeddings, the task encoder extracts the task-specific features corresponding to the special $<TASK>$ token. Finally, the task-specific predictor outputs the final answer from the task-specific features of the integrated multimodal data. In the following, we will describe each of these components in detail.
@@ 4.2.2 CTViT
CTViT extracts embedding features of multi-scale 3D CT volumes with physical size awareness. CTViT has two parts: a multi-scale CT tokenizer and an image encoder. To process a 3D CT scan, we divide each image volume into non-overlapped 3D patches as in [45]. Each 3D patch is referred to as an image token. Since various diseases are at different scales in the CT images, we design a multi-scale CT tokenizer, which consists of multiple linear embedding layers corresponding to different sizes of image patches, as shown in Figure 6-(a). Each embedding layer has a linear transformation and a set of learnable positional embeddings. Each image token embedding is the sum of its linear transformation and the positional embedding. All sizes of image tokens are mapped to the same image embedding space. Inspired by [41], we decompose the 3D position embedding into two parts indexing in-plane and through-plane positions respectively. In other words, we have two positional embeddings: one for the 2D space within each slice and the other for the 1D range of slice position. The 3D positional embedding is the sum of them. By doing so, the number of learned parameters can be reduced. Figure 6(c) shows the image encoder in detail. Different CT scans may have different physical sizes specific to the individual patient size. The physical size is an important factor in some clinical tasks. Thus, we encode the physical size for image tokens with sine-cosine functions of different frequencies and add it to the image token embedding. Then, the image encoder was implemented as the plain ViT [46] that consists of multiple self-attention Transformer layers and a subsequent linear transformation layer that maps the image embedding space to the task embedding space. By disentangling physical size from image contents, we can flexibly perceive any size of CT volumes with physical size awareness without resampling CT volumes to have a consistent image tensor across different inputs.
@@ 4.2.3 Text Transformer
Any decent language model can be utilized as the text Transformer in M3FM. Here the text Transformer was implemented as a Byte-level Byte-Pair-Encoding (BBPE) tokenizer [39], a text encoder consisting of the original Transformer layers [47], and a linear transformation layer, as shown in Figure 6-(d). On one hand, the text encoder encodes patient-specific clinical information, such as demographics, smoking history, disease history, cancer history, and other clinical data, which are free text; e.g., “The patient is 56.0 years old. Gender is Female. Ethnicity is neither Hispanic nor Latino. Height is 60.0 inches. Weight is 105.0 pounds. Education is associate degree/ some college. Former smoker. Smoking duration is 38.0 pack years. Smoking intensity is 20.0 cigarettes per day. 2.0 years since quit smoking. The patient had asthma (childhood) diagnosed at 7.0 years old. The patient had hypertension diagnosed at 53.0 years old. The patient had pneumonia diagnosed at 50.0 years old. Patient’s brother(s) (including half-brothers) have lung cancer.” On the other hand, the text encoder encodes free-text task instructions/questions, which are used as the input of the task encoder to extract task-specific embedding features from the multimodal data; e.g., “Is there any significant cardiovascular abnormality?” and “Predict the lung cancer risk over 6 years.” This approach allows for embedding any combination of clinical information through free-text prompting, regardless of order. The control signals for specific tasks are then extracted from the text prompts by the same text encoder. Again, the linear transform maps the text embedding space to the unified task embedding space. For all our M3FMs, the number of Transformer layers is 12, and the size of text token embeddings is 768.